The cat food is sold in tins containing 500 g.

Write down the matrix M such that the product XM will show, for each brand,

the total cost, in cents, of buying ten tins at both stores during May.


1
Expert's answer
2021-05-07T11:51:38-0400

"X=\begin{bmatrix}9.4&9.6\\10.2&10.4\\11.4&11.0\end{bmatrix}" , "Y=\begin{bmatrix}9.8&10.0\\10.2&10.2\\12.0&12.0\end{bmatrix}"


They represent the difference in price between May and June in each store for the 3 different brands.


Total cost for each brand "=\dfrac{500}{10}=50"


Product XM is-


"XM=\begin{bmatrix}9.4&9.6\\10.2&10.4\\11.4&11.0\end{bmatrix}\begin{bmatrix}50\\50\end{bmatrix}"


Ten The Matrix "M=\begin{bmatrix}50\\50\end{bmatrix}"